OpenStudy (rayford):

Select the best possible first step to solving the following system by first eliminating the x variable. 3x – 2y = –4 x – 5y = 16

OpenStudy (lgbasallote):

if you look at the x variables you have 3x and x. to eliminate x you need to multiply the second equation by -3 so it will look like 3x - 2y = 4 -3x + 15y = -48 now you can cancel x

OpenStudy (rayford):

ok im with you..

OpenStudy (rayford):

where did you go man :(]??

OpenStudy (lgbasallote):

hmm?

OpenStudy (lgbasallote):

reread my reply again;...i just mentioned what the best first step is :D

OpenStudy (rayford):

oh....well then....

OpenStudy (rayford):

*cough* good day sir..thank you

OpenStudy (lgbasallote):

haha welcome :D